sinxcosy =1/4 , 3tanx = 4tany , then sin (x-y)=?
a). 1/16 b). 1/8 c).3/16  d). 3/4

Grade:11

1 Answers

Akhil
6 years ago
3tan x=4tany,3sinx/cosx=4siny/cosy. 3sinx×cosy=4siny×cosx (sinx×cosy=1/4). 3×1/4=4×siny×cosxSiny×cosx=3/16Now,, Sin(x-y)=sinx×cosy-cosx×siny=1/4-3/16=1/16